7. (a) Use the rules of Boolean algebra provided in (5) to simplify the following logic



gate circuit:



(b)Draw the simplified logic gate circuit in (a)

1
Expert's answer
2021-11-25T16:05:55-0500

a)

for example:



output is:

"AB + BC(B + C)"


simplify:

"AB + BC(B + C)=AB+BBC+BCC=AB+BC+BC="

"=AB+BC=B(A+C)"


b)

simplified logic gate circuit: